x+y / x-y = 5/8 . x/y ?
Answer by stanbon(75887)

x+y / x-y = 5/8 . x/y ?
----
8(x+y) = 5(x-y)
------
8x-5x = -8y-5y
3x = -13y
x/y = -13/3
-----------------
Cheers,
Stan H.
---------